OpenStudy (anonymous):

Find the midpoint between the points (4,-3) and (-6,2).help

OpenStudy (ajprincess):

Let the coordinates of the midpoint be (x,y). \[x=\frac{ x_1+x_2 }{ 2 }\] \[=\frac{ 4-6}{ 2 }\] \[y=\frac{ y_1+y_2 }{ 2}\] \[=\frac{ -3+2 }{ 2}\]

OpenStudy (anonymous):

WELL MID-POINT FORMULA IS :